For what values of x is the graph of y=xe^−2x concave down?
1
Expert's answer
2020-03-10T12:48:44-0400

Graph of function f(x) is concave down when f"(x) < 0


"y=xe^{-2x}"

Differentiating with respect to x

"y'=e^{-2x}+x\cdot e^{-2x}(-2)"

"\Rightarrow y'=e^{-2x}\left(1-2x\right)"

Differentiating with respect to x again

"y''=e^{-2x}(0-2)+(1-2x)e^{-2x}\cdot(-2)"

"\Rightarrow y''=e^{-2x}(-2-2(1-2x))"

"\Rightarrow y''=e^{-2x}(-2-2+4x)"

"\Rightarrow y''=(4x-4)e^{-2x}"

point of inflection when "y''=0"

Since "e^{-2x}" is always positive

"(4x-4)=0"

"x=1"

We need to check sign of "y''" for x < 1 and x >1

When x < 1 ( let x = 0)

"y'' = e^{-2 \cdot 0}(4 \cdot0-4)=1 (0-4)=-4 < 0" ................concave downward

when x > 1 (let x=2)

"y''=e^{-2 \cdot 2}(4 \cdot 2-4)=e^{-4}\cdot 4 \approx 0.07 >0" ............concave upward


Thus, graph of "y=xe^{-2x}" is concave down for x <1

In interval notation, graph of "y=xe^{-2x}" is concave down in "(-\infty, 1)"